OpenStudy (anonymous):

Suppose f(π/3) =4 and f’(π/3)=-2, and let g(x)=f(x) sin x and h(x)= (cos x)/f(x). Find (a) g’(π/3) (b) h’(π/3)

jimthompson5910 (jim_thompson5910):

g(x)=f(x) sin x g'(x)=f'(x) sin(x)+ f(x) cos(x) ... use the product rule here g'(π/3)=f'(π/3) sin(π/3)+ f(π/3) cos(π/3) g'(π/3)=(-2)(sqrt(3)/2)+ (4)(1/2) g'(π/3)=-sqrt(3)+2 I'll let you do the other one

OpenStudy (anonymous):

how did you get the -2 and 4

jimthompson5910 (jim_thompson5910):

the -2 comes from f’(π/3)=-2 and the 4 comes from f(π/3) =4, both of which are given

OpenStudy (anonymous):

oops! my eyes must of just skip that, lol. Thanks!